OpenStudy (anonymous):

please help, make x the subject of the following: x+3=2a+bx

jimthompson5910 (jim_thompson5910):

\[\Large x+3 = 2a + bx\] \[\Large x-bx = 2a-3\] \[\Large x(1-b) = 2a-3\] \[\Large x = \frac{2a-3}{1-b}\]

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

the aim is to get all terms in x to one side and numbers to the other x + 3 - bx = 2a x- bx = 2a - 3 x(1 - b) = 2a-3 x = (2a-3) / (1 - b)